Subject: APO 3.5m users committee minutes, 9/14/09
From: Suzanne Hawley
Submitted: Wed, 23 Sep 2009 12:04:41 -0700 (PDT)
Message number: 1145
(previous: 1144,
next: 1146
up: Index)
****** in case you don't read this whole message, be aware
that the main action item is to send science highlights
to Suzanne Hawley, Mark Klaene and your Users committee
representative before Oct 5, 2009. All email addresses
are above. *********************************************
APO 3.5-m Users Committee Phonecon, 9/14/09
Attending: Suzanne Hawley, Ed Turner (for Michael Strauss), Scott
Anderson, Mark Klaene, Guy Stringfellow (for John Bally),Remy
Indebetouw, Bruce Gillespie
Absent: Jon Fulbright, Al Harper, Jon Holtzman
**********************************
User feedback, comments from institutional representatives:
Princeton (Ed Turner) - Ed had not heard of any user comments.
Colorado (Guy Stringfellow) - Guy had not heard of any user comments.
Johns Hopkins (Jon Fulbright) - No report.
Washington (Scott Anderson) - Scott said that a group of UW students
are descending on APO later this week.
Chicago (Al Harper) - No report.
New Mexico State (Jon Holtzman) - No report.
Virginia (Remy Indebetouw) - Remy mentioned the upcoming UVa student
trip to APO in November.
**********************************
Discussion of telescope/instruments report:
Regarding the more detailed report below, Mark said that it covers a two
month period, mostly in which the telescope was shut down for upgrades
and annual engineering maintenance activities. The primary mirror was
washed in situ again, with good results. July was mostly lost to bad
weather, while August and September were mainly spent on the
conversion to direct drives for altitude and azimuth. We now have
functioning direct drives on all three telescope motion axes, using
largely identical hardware and software. The installations went
fairly quickly, but the debugging and tuning of the axes controls took
more time. At this point, the new drive systems are declared to be
operational for science observations, but further tweaking and
characterization will continue over the next few months. Users should
be watchful for tracking and guiding anomalies, especially when their
targets cross the zenith or meridian, and during windy conditions--any
problems noted should be reported to the APO night and day staff.
Also, one of the operational benefits of the new drives is that the
primary mirror covers can now be opened and closed while tracking,
i.e., not only at the zenith position as with the old drives. This
will enable observers to obtain flat-field and wavelength calibrations
while simultaneously tracking their science targets.
There is not much to report on the instruments. Their annual
maintenance procedures were accomplished, new bigger ion pumps were
installed on DIS, Agile now has 2-inch filter holders and improved
filter-change capability. Agile still needs more scattered light
abatement (ongoing); users should still plan to use sky flats with
Agile. Bill Ketzeback will transition to the 3.5-m day staff crew,
and Nick McDonald is leaving later this year. Joe Huehnerhoff has
joined the staff to replace Bill as an Observing Specialist.
*************************
3.5-m Telescope, Instruments, and CIF Projects Highlights, 7/7/09
through 9/9/09
Mark Klaene
0) Overview
July saw a one-week shutdown for mirror cleaning and azimuth
controller software work. Some observing was scheduled with limited
results due to weather, as is typical during the summer months. August
and early September were devoted primarily to the direct-drive motor
installation on azimuth and altitude axes. To date, we have resumed
science observations with all summer work completed. However, we do
anticipate further commissioning and characterization work to continue
on the direct drives and new axis-controller software over the next
several months.
1) Telescope
Both azimuth and altitude were converted to direct-drive motors
(similar to the NA2 and TR2 rotators). These four axes also now run
with similar hardware and software controllers. The mechanical
installations went very well on azimuth but were a little more
involved for altitude--minor machining was required to get the proper
preload set for azimuth. The mechanical installation was 85% complete
after the first week and 100% complete after week two. Electrical
work went well. The most significant issue is that despite
manufacturer claims, the amplifiers can overheat in some conditions so
additional cooling was required. The remaining time consisted of
getting the axis controller software running with the new motors. This
was no small task and required a lot of tuning and re-configuring.
This is not 100% complete and we will continue fine-tuning over the
next several months.
We did not get all of our commissioning data completed due to weather,
but current performance does appear quite good.
2) Instruments
SPIcam, DIS, GFP were evacuated for normal annual servicing. DIS also
received two new ion pumps. The new pumps have three times the pumping
capacity as the old ones, which will provide some headroom for leaks
and permeation and also helps when servicing, allowing a quicker pump-
down time.
The echelle upper tank was evacuated. NIC-FPS and TripleSpec were kept
cold; no servicing work was accomplished on these.
Agile rotational pin and filter wheel were worked on. The 2" filter
holder is now in place. Work was done on the stray light seen in the
dome flats, however additional work is required.
3) CIF projects
These are essentially complete for this year with the new axis
controller, direct-drive motors and the Agile move to TR2.
4) Personnel
Bill Ketzeback will be moving over to day shift in the upcoming
months, he will be replacing Nick MacDonald as the 3.5-m Telescope
Engineer, who will be leaving APO at the end of the year.
Joe Huehenhoff has been hired to replace Bill on the night crew. He
will start training on the 10th of this momth.
**********************************
Instrument Studies (Hawley):
Suzanne gave an update on the ongoing new and upgraded instrument
studies. Regarding the echelle upgrade (led by Eric Burgh) and new
imaging camera (led by Cynthia Froning), Eric has been talking with
Dale Sandford regarding costs estimates for a new detector and optics
coatings. For the imaging camera, Jon Holtzman, Suzanne, and Cynthia
met with Chris Burrows to discuss a camera design for the widest
possible FOV at the NA2 instrument port. The e2v CCDs have smaller
pixels than SPIcam, so an f/2 optical system is required to yield a 10-
arcminute FOV. This would necessitate a major change in the NA2 light
baffling system. Another option is to look at an f/5 design that uses
a detector with larger pixels. This trade study should be completed
by the end of October. One of the wide-field design issues is that we
would need a new set of larger filters, and that a faster beam would
compromise the performance of narrow-band filters--we need to know how
important narrow-band filters are for users. Guy Stringfellow said
that the narrow-band filters are quite important. The f/5 design
might be made to work with our existing 3-inch filters. Another
option still being considered is the loan to APO of the QUOTA camera
from WIYN, which would entail possible significant requirements on
instrument control and user interface software.
We are also continuing to think about Jian Ge's FIRST instrument, for
which a white paper was distributed to the Users Committee last year.
FIRST is a NIR high-resolution spectrograph and interferometer that is
predicted to enable radial-velocity measurements with a precision of
~1 m/sec. A new wavelength calibration possibility using a laser
frequency comb developed at the University of Colorado is being considered
as part of this instrument. If members of our user community are interested
in single-target high-precision RV science programs in the near infrared
(e.g., exo-planet candidate follow-up, stellar oscillations, binaries,
etc.), they should express their interest to their Users Committee
representatives.
Suzanne added that since the direct-drives project is essentially
completed, we will increase the priority of instrumentation projects
in the next annual planning cycles.
Mark reported on the Goddard Fabry-Perot instrument. The GSFC team
has a new photon-counting detector for the GFP running in the lab back
in Greenbelt, MD. They also have the parts delivered for the IFU, and
the plan is to install the new detector and IFU in the GFP sometime in
the future. We may also place the GFP under TUI control, depending on
how wide the user interest is in the upgraded GFP. In the March/April
time frame, we posted a document that described the new performance
expectations for the upgraded GFP. Again, expressions of interest in
this instrument by users to their Users committee representatives
are desired.
**********************************
ARCSAT, SDSS-III status (Klaene):
ARCSAT (the 20-inch, aka "Photometric Telescope) has not been operated
recently. There are problems with its control system, and we are
getting in touch with DFM on diagnosing it. Since there hasn't been
much spare staff support time available during the shutdown season, we
hope to be able to get the telescope operational in the near future.
[note added 9/23: the problem was found and the telescope is now
operating nominally]
For SDSS-III, the imager is ready for its fall observing season, which
may be its last. MARVELS is operational, and the BOSS spectrograph
upgrades are installed and appear ready for commissioning tests.
There are issues being worked on guiding, operations software, and
plate marking and mapping. Lots of SDSS-III people have been on site,
and will continue to be so through the fall season as BOSS
commissioning is conducted.
**********************************
Science Highlights for Annual Report (Hawley):
Suzanne reminded the Users Committee members that it is the time to
provide their institutional science highlights for inclusion in the
Annual Report to the ARC Board of Governors. Rene Walterbos has
agreed to again gather and compile the science publications and theses
lists. Suzanne needs the science highlights--one page of graphics
plus text per highlight--submitted to Mark Klaene and herself not
later than 5 October. Multiple highlights from each institution are
encouraged. Users should send these to their Users committee
representatives as well as to Suzanne and Mark.
**********************************
ACTION ITEMS:
[open from previous months]:
none
[new actions from this meeting]:
===>ACTION: Suzanne needs the 2009 3.5-m telescope science highlights--
one page of graphics and text per highlight--submitted to Mark Klaene
and herself not later than 5 October. Multiple highlights from each
institution are encouraged, dating back to 1 October 2008.
**********************************
Next meeting: The next Users Committee phonecon will be on Monday, 5
October, at 8:30 AM Pacific Time. The agenda and other materials will
be sent to the committee members during the preceding week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 1145 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO